### v1

# Simple Yet Effective: Extracting Private Data Across Clients in Federated Fine-Tuning of Large Language Models

Abstract
Federated fine-tuning of large language models (FedLLMs) presents a promising approach for achieving strong model performance while preserving data privacy in sensitive domains. However, the inherent memorization ability of LLMs makes them vulnerable to training data extraction attacks. To investigate this risk, we introduce simple yet effective extraction attack algorithms specifically designed for FedLLMs. In contrast to prior "verbatim" extraction attacks, which assume access to fragments from all training data, our approach operates under a more realistic threat model, where the attacker only has access to a single client’s data and aims to extract previously unseen personally identifiable information (PII) from other clients. This requires leveraging contextual prefixes held by the attacker to generalize across clients. To evaluate the effectiveness of our approaches, we propose two rigorous metrics—coverage rate and efficiency—and extend a real-world legal dataset with PII annotations aligned with CPIS, GDPR, and CCPA standards, achieving 89.9% human-verified precision. Experimental results show that our method can extract up to 56.57% of victim-exclusive PII, with "Address," "Birthday," and "Name" being the most vulnerable categories. Our findings underscore the pressing need for robust defense strategies and contribute a new benchmark and evaluation framework for future research in privacy-preserving federated learning.
1 Introduction
Fine-tuning large language models (LLMs) using federated learning (FL) algorithms (Ye et al., 2024a, b; Zhang et al., 2023; Chen et al., 2024; Yao et al., 2024) enables distributed training across clients without requiring centralized data. Recent advances (Bai et al., 2024; Wu et al., 2025; Li et al., 2020; Karimireddy et al., 2020) have focused on algorithmic enhancements to improve performance and reduce training costs. However, the threat of training data extraction from federated LLMs remains largely underexplored.
LLMs are known to memorize and regurgitate portions of their training data (Carlini et al., 2021, 2023), including sensitive content such as Personally Identifiable Information (PII) (Shao et al., 2024; Kim et al., 2023; Nakka et al., 2024). In FL, clients retain local datasets and collaboratively train models by sharing intermediate updates (e.g., checkpoints) rather than raw data. Prior attacks often rely on unrealistic assumptions, such as the adversary possessing large portions of the training corpus (Yu et al., 2023; Huang et al., 2022).
In this work, we consider a more realistic threat model in which a malicious client exploits contextual prefixes from its own data to extract previously unseen PII from other clients. For example, a prefix like "The trial of this case has now concluded. The plaintiff, (Name)..." in the attacker client can be used to prompt the global model and elicit sensitive information from other clients. Herein, we propose three simple yet effective attack strategies: (1) collect all possible contextual prefixes from the attacker dataset for querying federated LLMs, (2) sampling high-frequency prefixes from the attacker’s dataset to query the federated LLM, and (3) locally fine-tuning on prefix–PII pairs in a supervised manner to increase extraction success rates.
We evaluate our attack methods in a realistic setting by annotating a real-world legal dataset with PII labels aligned with privacy regulations such as CPIS, GDPR, and CCPA (see Acronyms List A). To assess attack effectiveness, we define two metrics: i) coverage, the proportion of target PII from other clients that can be recovered, and ii) efficiency, the amount of PII extracted within a limited query budget.
The experiments on our dataset show that the proposed extraction attacks can achieve up to 56.57% coverage, with "Address," "Birthday," and "Name" identified as the most vulnerable PII types. However, we observe diminishing returns in coverage as the prefix budget increases. Our enhancement strategies improve the diversity of extracted PII under fixed budget constraints. These findings underscore the tangible privacy risks in FedLLMs and emphasize the need for further research into effective defense mechanisms. In summary, our contributions are three-fold:
  1. 1.
    We propose three simple yet effective training data extraction attack methods for FedLLMs. Their effectivess is assessed with two evaluation metrics: coverage and efficiency. The new attack approaches are orthogonal to the existing attacking methods, such as gradient reconstruction and membership inference.
  2. 2.
    We conduct comprehensive experiments demonstrating that these attacks can achieve up to 56.57% coverage of victim-exclusive PII, effectively extracting diverse PII types. Our results also reveal a trade-off between coverage and extraction efficiency.
  3. 3.
    We develop a real-world benchmark dataset by extending a legal corpus with PII annotations aligned with regulatory standards (CPIS, GDPR, CCPA), thus addressing the scarcity of high-quality datasets for PII extraction research in federated learning.
2 Related Work
This study is related to the fields of data extraction attacks and federated learning. For the reader’s convenience, a brief introduction to these concepts is provided in Section B. In this section, we only review the work that is directly related to our method.
2.1 PII Extraction Attacks in LLM
Large language models, due to their massive parameter scale, are capable of memorizing exact training data samples, making them vulnerable to data extraction attacks. These attacks can target different granularities of information: sample-level and entity-level.
At the sample level, an attacker with access to the full prefix of a training sample can query the LLM to regenerate the exact suffix (Yu et al., 2023; Shi et al., 2024; Zhang et al., 2024). This technique, known as verbatim training data extraction (Carlini et al., 2021, 2023; Schwarzschild et al., 2024), is widely used to detect data contamination and copyright violations (Dong et al., 2024).
At the entity level, attackers may know a subset of PII entities—such as names or affiliations—about a particular subject. By combining these known details with prompt templates (either manually crafted or automatically generated (Kassem et al., 2025)), they can elicit the model to produce additional PII records about the same subject. This is known as an associative data extraction attack (Shao et al., 2024; Kim et al., 2023; Zhou et al., 2024).
A PII extraction attack broadly refers to any attack aimed at eliciting outputs from the model that contain PII (Lukas et al., 2023; Nakka et al., 2024; Huang et al., 2022). Both verbatim and associative techniques can be used to conduct such attacks.
2.2 Privacy Threats in Federated Learning
Threats in Federated Learning can be categorized into two main areas: security and privacy(Wang et al., 2024; Xie et al., 2024; Li et al., 2024a). Security threats typically aim to disrupt the entire FL system by invalidating model training(Shejwalkar and Houmansadr, 2021) and introducing backdoors(Bagdasaryan et al., 2020; Chang et al., 2024). In contrast, privacy threats have attracted more attention from researchers and focus on stealing confidential information from the FL system, such as inferring sensitive properties(Melis et al., 2019), reconstructing clients’ private datasets(Zhu et al., 2019; Geiping et al., 2020), and determining the membership and source of training data(Rashid et al., 2025; Vu et al., 2024; Hu et al., 2024). To achieve these attacks, researchers often make different assumptions regarding the attacker’s knowledge. Common assumptions typically fall into two dimensions: whether the attacker is a client or a serverChu et al. (2023), and whether the attacker is semi-honest(Applebaum, 2017; Hu et al., 2024) or malicious. These assumptions determine whether the attacker has access to gradients, local datasets, model parameters and the ability to manipulate them.
3 Dataset
3.1 Data Sources and Preprocessing
The majority of our dataset comes from the Challenge of AI in Law (CAIL) (Li et al., 2024b), supplemented by smaller portions from CJRC (Duan et al., 2019) and JEC-QA (Zhong et al., 2020). CAIL is a renowned annual competition offering a variety of tasks in legal NLP. We specifically focus on two natural language generation (NLG) tasks-SUM and RC-and three NLP tasks-Match, Exam, and Cls. Detailed descriptions of each task can be found in Section D and Table 4 lists examples of each task.
Prior work (Zhang et al., 2023; Yue et al., 2023) has employed these datasets for LLM and FedLLM research. Based on their outputs, we further preprocess and curate the data, as described in Section E. Table 5 provides detailed dataset statistics.
3.2 PII Labeling
We reviewed the definitions and examples of PII in various legal provisions, including CPIS, GDPR, CCPA and Singapore PDPC (see Acronyms List A), and used them as references to establish a systematic PII labeling standard. We selected PII types relevant to the text modality, removing those related to the biological modality (e.g., facial recognition, fingerprint, DNA). We also removed PII types that are unlikely to appear in legal texts (e.g., browser history, SMS content, IP & MAC addresses), as well as those that are difficult to describe or evaluate (e.g., medical examination reports, psychological trends). Ultimately, we defined labeling guidelines encompassing 7 major categories and 36 subcategories. These standards are summarized in Table 6, and the distribution of labeled PII types is shown in Figure 1.
Figure 1: Distribution of de-duplicated PII instances by label category.
We employed a combination of machine-assisted annotation and manual verification to label the data. For each major category of PII, we designed a specific prompt (as shown in Figure 5) and utilized GPT-4o (OpenAI et al., 2024) for initial labeling. Subsequently, we used Label Studio (Tkachenko et al., 2020-2025) and recruited five students with a foundational understanding of legal concepts and judgment skills to evaluate a subset of the annotated data. We selected 100 samples, which included over 6,000 PII labeling records, and evaluated them. The results of this evaluation are presented in Table 7.
4 Method
4.1 Problem Definition
We investigate a novel class of data extraction attacks that is particularly realistic in the emerging federated large language model (FedLLM) setting. This attack fundamentally differs from traditional verbatim data extraction in its assumptions, setup, and goals.
First, unlike prior work where the attacker is assumed to have access to most of the training data, our setting assumes that the attacker possesses only a small subset of the overall training corpus. Second, in our formulation, the prefix and target suffix do not form a contiguous span within the original training data. Specifically, extraction prefixes are sampled from the attacker’s local dataset $D_{a}$, while the target suffixes exist solely in other clients’ data and are absent from $D_{a}$. As a result, each prefix must generalize across contexts to trigger the generation of previously unseen suffixes. Third, the attacker is not interested in extracting all training completions, but rather focuses on recovering specific, sensitive content, such as personally identifiable information (PII).
Formally, we consider an FL system comprising $c$ clients $\mathcal{C}=\{C_{1},C_{2},\dots,C_{c}\}$, where each client $C_{i}$ holds a local dataset $D_{i}$. Among them, one client—denoted as the attacker $C_{a}\in\mathcal{C}$—is assumed to be semi-honest (Applebaum, 2017; Hu et al., 2024). That is, $C_{a}$ faithfully follows the FL protocol (e.g., does not poison data or manipulate model weights), but acts adversarially in a passive manner, attempting to infer PII contained in other clients’ datasets by analyzing the global model $\theta$.
In this setting, the attacker issues queries to the model $\theta$ to extract data without knowing which client any particular output originates from. However, for evaluation purposes, we designate one client as the reference victim to measure the attack’s effectiveness. Let $S_{a}$ and $S_{v}$ denote the sets of PII instances held by the attacker and the victim client, respectively. The attacker constructs a prompt set $P$ and queries the federated language model $\theta$, obtaining a corresponding output set $Y$. We formalize key definitions and evaluation metrics in the following subsections.
Definition 1 (Extracted).
A PII instance $s\in S_{v}$ is considered successfully extracted if there exists a prompt $p\in P$ and a corresponding model output $y\in Y$ such that:
$\exists u\in\Sigma^{*}\text{ such that }y=s\oplus u,$ (1)
where $\Sigma^{*}$ is the set of all finite-length strings over the vocabulary, and $\oplus$ denotes string concatenation. In other words, the model output $y$ begins with $s$.
Definition 2 (Coverage Rate (CR)).
The coverage rate measures how thoroughly the attacker recovers the PII unique to the victim client. It is defined as:
$S_{E}$ =$=\{s_{i}\mid\exists y\in Y\text{ such that }s_{i}\text{ is extracted by }y\},$
CR $=\frac{|(S_{v}\setminus S_{a})\cap S_{E}|}{|S_{v}\setminus S_{a}|}.$ (2)
A higher CR indicates that a larger fraction of the victim’s unique PII has been successfully extracted.
Definition 3 (Efficiency (EF)).
Efficiency quantifies the precision of extraction with respect to the number of queries. Let $Q$ denote the number of queries, the efficiency is defined as:
$\text{EF}=\frac{|(S_{v}\setminus S_{a})\cap S_{E}|}{Q}.$ (3)
A higher EF indicates that more PII is extracted with fewer queries.
Building upon these definitions, the central challenge is to design algorithms that enables the attacker to extract PII both comprehensively and efficiently—that is, achieving high coverage (CR) and high efficiency (EF).
4.2 Attacking Algorithms
4.2.1 PII-contextual Prefix Sampling
We begin with a straightforward method for constructing query prompts using PII-contextual prefixes—text segments that immediately precede PII instances in the attacker’s own dataset $D_{i}$. This approach is motivated by the observation that manually crafted prompts such as "my phone number is" may not align with the model’s training distribution, and are often ineffective at extracting attacker-defined PII instances.
Let the attacker’s training corpus be denoted as $U_{a}=t_{0}t_{1}\cdots t_{|U_{a}|}$, formed by concatenating all samples in $D_{i}$, where each $t_{i}$ represents a word token. Let $\mathcal{S}$ be the multiset of PII instances labeled in $U_{a}$. Define the function $\text{Loc}(s)$ as the index of the first word of a PII instance $s$ in $U_{a}$. We further define a contextual prefix extraction function $\mathcal{T}_{\lambda}(U,s)$ that returns the $\lambda$-length prefix ending just before $s$:
$\mathcal{T}_{\lambda}(U,s)=t_{\text{Loc}(s)-\lambda}\cdots t_{\text{Loc}(s)-1}.$
The resulting set of PII-contextual prompts is given by:
$P_{c}=\{\mathcal{T}_{\lambda}(U_{a},s)\mid s\in\mathcal{S}\}.$ (4)
Once $P_{c}$ is constructed, the attacker $\mathcal{C}_{a}$ uses the global model $\theta$ to generate candidate PII completions. For each prefix $p\in P_{c}$, a suffix $y$ of at most $m$ tokens is sampled according to the model’s conditional distribution:
$y=x_{1}x_{2}\dots x_{m}\sim\mathbf{P}(y\mid p;\theta).$
To increase the diversity of generations, $\mathcal{C}_{a}$ may sample $n$ independent suffixes for each prefix $p$, forming:
$Y_{p}=\{y_{1},y_{2},\dots,y_{n}\}.$
The final set of candidate generations is the union across all prefixes:
$Y=\bigcup_{p\in P_{c}}Y_{p}.$
This results in a total query cost of $Q=n\cdot|P_{c}|$.
We can further generalize the contextual prompt set $P_{c}$ by collecting a multiset of all prefix substrings that end immediately before each PII instance:
$\text{SUP}(P_{c})=\{t_{i}\cdots t_{\text{Loc}(s)-1}\mid(\text{Loc}(s)-i)\in[1,\lambda]\}.$
This produces a much larger set of contextual prefixes, which may yield a higher extraction coverage rate (CR) but an extremely low efficiency (EF) due to the massive number of query prompts.
4.2.2 Frequency-Prioritized Prefix Sampling (FP Sampling)
We hypothesize that prefixes appearing more frequently immediately before PII entities are more strongly associated with different PII instances. Based on this intuition, our goal is to prioritize such high-frequency prefixes to construct a more compact and informative prefix set.
To formalize this, we partition the multiset $\text{SUP}(P_{c})$ based on prefix frequency. For each integer $\sigma\geq 1$, let:
$P_{\sigma}=\{p\in\text{SUP}(P_{c})\mid\text{Count}_{\text{SUP}(P_{c})}(p)=\sigma\}.$
which defines subsets of prefixes that occur exactly $\sigma$ times. This induces a frequency-based decomposition of the unique prefixes in $\text{SUP}(P_{c})$:
$\text{Set}(\text{SUP}(P_{c}))=\bigcup_{\sigma\geq 1}P_{\sigma}.$
Given a frequency threshold $\sigma_{a}$, we define the set of frequent prefixes as:
$P_{f\geq\sigma_{a}}=\bigcup_{\sigma\geq\sigma_{a}}P_{\sigma},$ (5)
which is then sorted in descending order of frequency. Notably, setting $\sigma_{a}=1$ yields the full set of generalized contextual prefixes, $P_{f\geq 1}=\text{Set}(\text{SUP}(P_{c}))$, sorted in descending order of frequency.
Given a prefix budget $B$, we construct the final prefix set by selecting the top-$B$ prefixes from $P_{f\geq\sigma_{a}}$, thereby prioritizing high-frequency prefixes during sampling.
4.2.3 Latent Association Fine-tuning (LAFt)
We hypothesize that a model’s susceptibility to extraction attacks stems from its latent ability to associate two underlying conceptual distributions: (1) $\mathcal{A}$ — the distribution over prefixes that are semantically or syntactically likely to precede PII; and (2) $\mathcal{B}$ — the distribution over actual PII instances. Let $\text{Dist}(\mathcal{A},\mathcal{B};\theta)$ denote the degree of association between these two distributions under model parameters $\theta$.
Because this association is implicitly encoded in the model’s internal representations, we propose a method to reduce this distance through parameter updates—a technique we term Latent Association Fine-tuning (LAFt). The core idea is to fine-tune the model to minimize $\text{Dist}(\mathcal{A},\mathcal{B};\theta)$, thereby reinforcing its internal linkage between indicative prefixes and corresponding PII, ultimately improving its capacity for PII extraction.
To implement this, we construct a fine-tuning dataset $D_{\text{ft}}$ by pairing frequent prefixes with known PII instances from the attacker’s dataset:
$D_{\text{ft}}=\{(p,s)\mid p\in P_{f},\ s\in S_{a}\},$ (6)
where $P_{f}$ is the set of frequent prefixes derived from $D_{a}$, and $S_{a}$ is the attacker’s known PII set. We then fine-tune the model using the standard causal language modeling objective:
$\theta^{\prime}=\arg\min_{\theta}\sum_{(p,s)\in D_{\text{ft}}}\sum_{t=1}^{|s|}-\log\mathbf{P}(s_{t}\mid p,s_{<t};\,\theta).$
The updated model $\theta^{\prime}$ is then used for extraction, using prefixes from either $P_{f}$ or $P_{c}$.
5 Experiment
5.1 Experimental Setup
We detail the experimental setup and implementation to ensure reproducibility.
Federated Setup.
We simulate a federated system with 5 clients using a label-skewed non-IID data partitioning method based on clustering of language embeddings (Li et al., 2023). Each client receives a comparable number of samples. Federated fine-tuning is conducted on legal tasks using the OpenFedLLM framework Ye et al. (2024b), with FedAVG as the aggregation method over 10 communication rounds. All clients adopt parameter-efficient fine-tuning (LoRA) and a shared prompt template. Hyperparameter settings and implementation details are detailed in Appendix H.
After federated fine-tuning, we evaluate the utility of the final global model on a held-out global test set. Following common practice, we compare this with a centrally trained (non-FL) baseline on the same test set. Results are shown in Table 8.
Models and Metrics.
We use QWen1-8B (Bai et al., 2023) and Baichuan2-7B (Yang et al., 2023), both pre-trained primarily on Chinese corpora. Model performance is evaluated with Coverage Rate (CR), Efficiency (EF), and Victim-exclusive Extracted PII (VxPII, defined as $|(S_{v}\setminus S_{a})\cap S_{E}|$), defined to measure extraction accuracy and completeness.
Attack Strategies.
We designate client 0 as the attacker and client 1 as the victim, and evaluate three strategies: (1) PII-contextual prefix sampling. The attacker builds a prefix set $P_{c}$ from its local dataset $D_{0}$ with prefix length $\lambda=50$. Each prefix queries the global model 15 times, generating up to $m=10$ tokens per query—sufficient to recover most PIIs with manageable cost. (2) Frequency-prioritized sampling. Prefixes in $\text{Set}(\text{SUP}(P_{c}))$ are ranked by frequency to form $P_{f\geq 1}$ and used in descending order. Sweeping the prefix budget $B$ (base-10 exponential scale) varies the frequency threshold $\sigma_{a}$, enabling analysis of coverage–efficiency trade-offs. (3) Latent association fine-tuning. The attacker fine-tunes the global model (1 epoch, LR = 5e-5, LoRA: $r=16$, $\alpha=32$) using 10k frequent prefixes and 10k randomly sampled PIIs from its own data to reinforce prefix–PII associations.
Further implementation details are provided in Appendix H.3.2.
Evaluation Protocol.
To ensure fair evaluation, we define the set of victim-exclusive PIIs as $(S_{v}\setminus S_{a})$ and apply two filters: (1) only PIIs $s_{i}\in S_{v}$ that do not appear in the attacker’s training corpus ($s_{i}\notin U_{a}$) are retained; and (2) we enforce a longest common prefix constraint to eliminate interference between PIIs—e.g., those sharing prefixes—which may confuse the determination of which PII has been extracted (see Equation (1)):
$\text{LCP}(s_{i},s_{j})=0,\quad\forall s_{i}\neq s_{j}\in S_{v}$
Metrics are computed on this filtered and prefix-disjoint set, as defined in Equations (2)–(3).
5.2 Main Results
5.2.1 RQ1: To what extent can such attacks extract PII at maximum?
We begin by evaluating the maximum coverage rate (CR) achievable by our proposed attack methods. Specifically, we use the full set of PII-contextual prefixes $P_{c}$ to query the federated fine-tuned LLMs. To explore the upper bound of extraction capability, we also employ the sorted generalized contextual prefix set, $\text{Set}(\text{SUP}(P_{c}))=P_{f\geq 1}$, though this is limited to the Qwen1-8B model due to its high computational cost.
Table 1 reports the results. Using the PII-contextual prefix set $P_{c}$, we achieve a CR of 22.93% on Qwen1-8B and 28.95% on Baichuan2-7B. Applying Latent Association Fine-tuning (LAFt) improves Qwen1-8B’s CR to 28.30%, though the corresponding efficiency (EF) remains low at 0.24%, indicating a high proportion of irrelevant or duplicated outputs. The count of unique, victim-exclusive PII (VxPII) remains relatively small.
Notably, leveraging the broader prefix set $P_{f\geq 1}$ substantially increases Qwen1-8B’s CR to 56.57%. However, this comes at the expense of extremely low efficiency (0.01%), suggesting that while coverage increases, most samples are not useful.
For Baichuan2-7B, applying LAFt does not lead to a direct improvement in CR, but we later show that LAFt consistently recovers a distinct set of PII instances.
Table 1: Attack performance of client 0 (attacker) against client 1 (victim) using different methods. The $(S_{v}\setminus S_{a})$ contains 8870 unique victim-exclusive PII items.
Model Prefix Set LAFt CR EF VxPII Count Prefix Set Size
Qwen1-8B $P_{c}$ without 22.93% 0.1910% 2034 71006
$P_{c}$ with 28.30% 0.2357% 2510 71006
$P_{f\geq 1}$ without 56.20% 0.0110% 4985 3013161
$P_{f\geq 1}$ with 56.57% 0.0111% 5018 3013161
Baichuan2-7B $P_{c}$ without 28.95% 0.2411% 2568 71006
$P_{c}$ with 28.46% 0.2370% 2524 71006
5.2.2 RQ2: What is the distribution of extracted PII?
Figure 2 presents the label distribution of deduplicated victim-exclusive PII instances extracted by the Qwen1-8B and Baichuan2-7B models. The horizontal bar plot displays label names on the y-axis and the corresponding instance counts on the x-axis.
Figure 2: Label distribution of deduplicated victim-exclusive PII instances.
Among all extracted PII types, "Address", "Birthday", and "Name" appear most frequently. In contrast, categories such as "Work Experience" and "Work Place" occur less often but remain notable. More complex categories, such as "Medication Record", are not extracted at all. This absence is primarily due to the nature of the extraction attack and the evaluation protocol, which only considers model-generated PII that exactly matches the ground truth. Since complex PII types often manifest as long free-text segments, it is difficult for models to reproduce them with verbatim accuracy.
5.2.3 RQ3: How Effective Are FP Sampling and LAFt in Extracting VxPII?
We present a Venn diagram of the VxPII (Victim-exclusive extracted PII) sets produced by different methods in Figure 3. As shown, the Frequency-Prioritized Prefix Sampling (FP Sampling) and the Latent Association Fine-tuning (LAFt) methods identify overlapping but distinct sets of PIIs, , suggesting that each method captures different aspects of information.
We also analyze the VxPII distribution produced by the LAFted model and find that, although LAFt uncovers a different subset of instances, the overall label distribution remains similar across models. This indicates that LAFt tends to uncover new instances of already-well extracted PII types, rather than broadening the diversity of PII types. This phenomenon may be attributed to the way we construct the LAFt fine-tuning dataset $D_{\text{ft}}$, where frequent prefixes are paired with random PIIs. To encourage the discovery of rarer PII types, future work might explore pairing prefixes with less frequent or underrepresented PII types during fine-tuning.
Figure 3: Venn diagrams showing the overlap of VxPII sets extracted by different methods. (a) Comparison of PII-contextual prefix set $P_{c}$ vs. frequency-prioritized prefix set $P_{f\geq 1}$ with a prefix budget of $B=10000$, on Qwen1-8B without LAFt; (b) Comparison of VxPII extracted by Qwen1-8B with and without LAFt, using the full $P_{c}$ prefix set; (c) Comparison of VxPII extracted by Baichuan2-7B with and without LAFt, using the full $P_{c}$ prefix set.
5.2.4 RQ4: What Is the Tradeoff Between Coverage and Efficiency?
Figure 4 illustrates the tradeoff between coverage rate (CR) and efficiency (EF) as the prefix budget $B$ increases. Specifically, CR rises significantly while EF declines sharply for both prefix sets $P_{c}$ and $P_{f\geq 1}$. As more PII are extracted, discovering new, non-duplicate PII becomes increasingly difficult. This observation points to a potential tradeoff decision in real-world attack scenarios, where expanding the budget could lead to diminishing efficiency gains.
Figure 4: Coverage rate (CR) and efficiency (EF) under varying prefix budgets $B$ for prefix sets $P_{c}$ and $P_{f\geq 1}$. The budget values are scaled exponentially (base 10).
5.3 Ablation Study
To evaluate the robustness of our PII extraction approach, we conduct a cross-validation experiment by iteratively designating each client as the attacker while treating the remaining clients as victims. This helps ensure that our method is not biased towards a particular client. As shown in Table 9, the proposed extraction attack demonstrates consistent effectiveness across all attacker–victim pairings.
5.4 Evaluating the Effectiveness of Simple PII Masking
We evaluate the effectiveness of a simple data sanitization strategy that masks PII using existing annotations. Each PII instance in the training data is replaced with a string of asterisks (*) of equal length. We then re-fine-tune FedLLM on this sanitized dataset and perform the PII-contextual Prefix Sampling attack once again. Table 2 compares the attack performance with and without the PII masking defense.
Table 2: Attack performance with and without PII masking using the contextual prefix set $P_{c}$. The model is Qwen1-8B.
VxPII CR EF
With PII Masking 2017 22.74% 0.1894%
Without Defense 2034 22.93% 0.1910%
As shown in the results, the number of extracted VxPII is only slightly reduced. To understand this, we examine the document frequency of VxPII instances—that is, how often each appears in the training data. Figure 7 shows that PII masking significantly lowers the frequency of most VxPII, suggesting that our annotations effectively cover the majority of PII. Interestingly, some VxPII with zero document frequency—indicating they were absent from the masked dataset—were still extracted.
Based on these observations, we identify two main reasons for the limited effectiveness of the masking defense: (1) Pretraining data contamination: Our training data, sourced from publicly available legal documents, likely overlaps with the pretraining corpora of models like Qwen1-8B and Baichuan2-7B. (2) Incomplete PII labeling: Some PII instances may remain unlabeled—and therefore unmasked—in the training data. In the realistic scenarios, attackers can customize PII definitions, making exhaustive and comprehensive coverage inherently challenging.
Pretraining data contamination is difficult to eliminate, as LLM providers rarely disclose their pretraining corpora. To mitigate this, we adopt a conservative strategy: we compare the VxPII extracted from FedLLM ($F$) with that from its un-fine-tuned base model ($B$), and subtract $B$ to isolate PII memorized during federated fine-tuning. Table 3 shows that even after removing $B$, a substantial number of VxPII remain in $F\setminus B$, confirming memorization during fine-tuning. Figure 8 further shows that $F\setminus B$ exhibits a similar label distribution of VxPII as observed in Figure 2, supporting the validity of our method.
Table 3: Comparison of VxPII sets between attacks on FedLLM and its un-fine-tuned base model.
Prefix Set Model $|F\setminus B|$ $|B\setminus F|$ $|F\cap B|$
$P_{c}$ Qwen1 682 518 1801
$P_{f\geq 1}$ Qwen1 554 308 4611
$P_{c}$ Baichuan2 407 405 2161
6 Conclusion
To investigate the privacy risks of data extraction attacks in realistic settings, we introduce a new class of attacks targeting FedLLMs. We extend a legal dataset with systematic PII annotations aligned with major privacy regulations, and evaluate attack performance using two key metrics: coverage rate and efficiency. Extensive experiments demonstrate that certain PII types are highly vulnerable, and our proposed methods can achieve substantial extraction performance. These findings highlight a critical privacy gap in FedLLMs and underscore the urgent need for stronger defense mechanisms in future federated learning systems.
Appendix A Acronyms List
  • GDPR - General Data Protection Regulation [European Union, 2016]
  • CCPA - California Consumer Privacy Act [State of California, US, 2018]
  • CPIS - Chinese Information Security Technology: Personal Information Security Specification (GB/T 35273-2020) [Standardization Administration of China (2020), SAC]
  • Singapore PDPC - Personal Data Protection Commission (Singapore) [Personal Data Protection Commission, Singapore, 2012]
  • Non-IID - Non-independent and identically distributed
Appendix B Preliminary
B.1 Data Extraction Attack
Early research on training data extraction attacks has broadly categorized them into untargeted and targeted attacks [Research, 2022, Yu et al., 2023]. Untargeted extraction aims to recover any memorized training samples without specifying a target [Lukas et al., 2023], whereas targeted extraction attempts to reconstruct specific training samples, often by providing a known prefix and recovering the remaining content [Carlini et al., 2021]. The latter type, often referred to as Verbatim Data Extraction, has become a standard approach for evaluating memorization in LLMs [Carlini et al., 2023, Dong et al., 2024] and for detecting potential data contamination [Dong et al., 2024]. We briefly outline the core methodology of verbatim data extraction below.
Given an LLM $\theta$ and a training dataset $X$, each training sample $x_{i}\in X$ is partitioned into two segments: a prefix $a_{i}$ and a suffix $b_{i}$, such that $x_{i}=a_{i}b_{i}$. The model is then prompted with $a_{i}$ to generate a completion $g_{i}$ of the same length as $b_{i}$. If $g_{i}$ exactly matches $b_{i}$, the sample is considered successfully extracted.
In practice, model outputs may not exactly replicate the original suffix but can still be lexically close. To accommodate this, a similarity-based metric such as Edit Distance [Levenshtein, 1965] is often employed. A sample is deemed extracted if the similarity score between $g_{i}$ and $b_{i}$ exceeds a predefined threshold $t$. By computing this similarity-based extraction score across all samples in a dataset $D$, one can quantify the model’s memorization behavior or assess its vulnerability to training data extraction attacks.
B.2 Federated Learning
Federated Learning (FL) is a solution to address data isolation issues [Yang et al., 2019], where a central server and multiple clients collaborate to complete the training process. A key feature of FL is that the training datasets are stored locally on each client and remain invisible to other clients. FL is commonly used in industrial scenarios where each client represents an independent organization, such as hospitals collaborating to train a medical model without combining their datasets due to legal restrictions or business competition. Federated Learning enables the training of stronger models compared to training on data from a single client alone.
Given $c$ clients and their private datasets $D_{1},D_{2},\dots,D_{c}$, the federated learning process aims to learn a global model $\theta$ by solving the following optimization problem:
$\theta^{*}=\mathop{\arg\min}\limits_{\theta}\frac{1}{c}\sum_{i=1}^{c}\mathcal{L}(D_{i},\theta)$
To solve this problem, many federated optimization algorithms have been proposed, such as FedAVG [McMahan et al., 2023] and FedProx [Li et al., 2020]. Typically, these algorithms consist of two alternating phases: local updating and central aggregation. In the local updating phase, each client independently optimizes the global model using its own dataset. In the central aggregation phase, the server aggregates the models from the clients using an aggregation algorithm, obtaining a global model, which is then sent back to each client for the next round of local updating. A typical procedure of federated learning is illustrated in Algorithm 1.
Appendix C A Framework of Federated Learning Algorithms
Algorithm 1 A Federated Learning Framework
1:Input: Clients set $\mathcal{C}=\{c_{1},c_{2},\dots,c_{c}\}$ with local datasets $D_{1},D_{2},\dots,D_{c}$; total FL rounds $R$; initial global model $\theta_{0}$; server aggregation function $f_{\text{agg}}$; client loss function $\mathcal{L}$
2:Output: Learned global model $\theta_{R}$
3:
4:ServerExecute:
5:for round $r=1$ to $R$  do
6:     for each client $c_{i}\in\mathcal{C}$ (in parallel) do
7:         $\theta_{r}^{i}\leftarrow$ ClientUpdate($c_{i},\theta_{r-1}$)
8:     end for
9:     $\theta_{r}\leftarrow f_{\text{agg}}(\{\theta_{r}^{i}|c_{i}\in\mathcal{C}\})$
10:end for
11:
12:ClientExecute:
13:function ClientUpdate($c_{i},\theta_{r-1}$)
14:     $\theta_{r}^{i}\leftarrow\mathop{\arg\min}\limits_{\theta}\mathcal{L}(\theta_{r-1},D_{i})$
15:     return $\theta_{r}^{i}$
16:end function
Appendix D Task Descriptions and Examples
  1. 1.
    Judicial Summarization (Sum): The task of judicial summarization aims to extract key information from court judgments and generate concise summaries. The input to this task is a legal document, and the output is a summary of its content. The performance of this task is evaluated using the Rouge-L metric, which effectively measures the similarity between the generated and reference texts based on the longest common subsequence (LCS). Rouge-L is a widely used metric in text generation tasks. In this study, we adopt Rouge-L because it captures both semantic and structural similarities between texts, making it suitable for summarizing judicial documents.
  2. 2.
    Judicial Reading Comprehension (RC): This task focuses on answering legal questions based on court documents to evaluate the model’s reading comprehension ability. The input consists of a piece of legal material and a question, and the task requires answering the question based on the content of the material. The performance metric for this task is Rouge-L.
  3. 3.
    Similar Case Matching (Match): In this task, the input includes three case documents, and the model is required to determine which of the latter two documents is more similar to the first one. The model selects the most similar document by computing the similarity between the first case and each of the other two. The evaluation metric for this task is accuracy.
  4. 4.
    Judicial Exam (Exam): This task simulates multiple-choice questions from legal examinations to assess the model’s knowledge of legal concepts. Given a judicial exam question with multiple options, the model is expected to choose the correct answer. The performance is evaluated using accuracy.
  5. 5.
    Legal Case Classification (Cls): This task requires the model to classify the cause of action in a case, assisting legal retrieval systems in automatically categorizing case types. The input is a description of the case facts, and the model is required to output the corresponding case category. The performance metric is accuracy.
Table 4: Input and Output Examples for Each Task
Task Input Output
Judicial Summarization
First-instance civil judgment on inheritance dispute between Han and Su
Shenyang Dadong District People’s Court
Plaintiff: Han, female, born June 6, 1927, Han ethnicity…
Clerk: Li Dan
Summary: This case involves an inheritance
dispute between the plaintiff and the defendant.
The plaintiff requests…
Judicial Reading Comprehension
Case: Upon trial, it was found that on February 11, 2014, the plaintiff…
Question: When did the plaintiff and defendant agree on the travel plan?
The plaintiff and defendant agreed on the
travel plan on February 11, 2014.
Similar Case Matching
Determine whether Case A is more similar to Case B or Case C.
A: Plaintiff: Zhou Henghai, male, born October 17, 1951…
B: Plaintiff: Huang Weiguo, male, Han ethnicity, resident of Zhoushan City…
C: Plaintiff: Zhang Huaibin, male, resident of Suzhou City, Anhui Province, Han ethnicity…
B
Judicial Exam
Wu was lawfully pursued by A and B… Which of the following analyses is correct?
A. If Wu missed both A and B, and the bullet…
B. If Wu hit A, resulting in A’s death…
C. If Wu hit both A and B, causing A’s death and B’s serious injury…
D. If Wu hit both A and B, causing both to die…
A
Legal Case Classification
Legal document: Plaintiff Yan Qiang submitted the following claims to this court:… Private Loan Dispute
Appendix E Data Preprocessing
Previous works [Zhang et al., 2023, Yue et al., 2023] have used these datasets for LLM and FedLLM research. In this work, we use the processed datasets from these prior studies and further curate the data for our experiments. We applied the following preprocessing steps to prepare the datasets:
Deduplication and Cleansing.
To ensure the quality of our data, we remove duplicate samples with logically equivalent meanings. For example, in the RC tasks, some samples only differ in the order of two legal cases. We also clean out samples containing garbled characters or large segments with a mixture of multiple languages.
Unifying Prompt Template and Instruction Reshaping.
Some tasks, such as Exam, contain instructions that appear in different parts of the sample (either at the beginning or the end). To standardize the format, we reshape the data so that the instruction always appears at the beginning, followed by the legal document. Additionally, we employ hierarchical hyper markers such as "<Case A>", "<Case B>", and "<Answer>" to clearly segment the prompt, making the structure more transparent for the LLM.
Appendix F Details of Dataset
Table 5: Dataset Statistics
Exam RC SUM Match Cls
#Samples 2399 3500 2651 3848 4196
Appendix G Details about PII Machine Annotation
Table 6: Categorization of Personally Identifiable Information (PII) Types in Our Labeling Standards
Major Category Minor Category
Personal Basic Information Name, Birthday, Address, Gender, Ethnicity, Family Relationship, Age, Nationality, Personal Phone Number
Personal Identity Information ID Number, Social Security Number, Driver’s License Number, Employee Number, License Plate Number
Health Related Information Physical Condition, Fertility Information, Current Medical History, Diagnosis and Treatment Status, Other Medication Record
Work and Education Information Workplace, Position, Work Experience, Education Experience, Grades
Personal Property Information Bank Account, Amount of Funds, Fund Flow Records, Virtual Assets, Other Financial Records
Personal Location Information Precise Location, Accommodation Information, Travel Trajectory
Others Marital History, Religious or Philosophical Beliefs, Sexual Orientation or Sex Life, Unpublished Criminal Records
I would like you to assist in reviewing the provided document and labeling all sections containing {{Major Categories of PII}} according to the following requirements.
1. **Types of personal information to identify include:**
{{PII Subcategories}}
2. **Output format:**
{{Output Format Description}}
3. **Input instructions:**
{{Input Format Description}}
Please provide the output directly in accordance with the format requirements above, without any additional explanation or comments. Thank you for your assistance!
Figure 5: PII Machine Annotation Prompt Template
Table 7: Human Evaluation of PII Labeling Quality
Evaluation Criteria Precision (P) Recall (R) F1 Score (F1)
Identical Span Only 0.89 0.93 0.91
Identical Span and Label 0.89 0.90 0.89
Appendix H Implementation Details
H.1 Federated Dataset Partitioning.
We use the preprocessed and labeled datasets (see Section 3) for our experiments, splitting the data into training and testing sets. In the federated learning setup, we simulate a system with five clients. The testing set remains global, while the training set is heterogeneously partitioned across the clients using a balanced Non-IID distribution (see Acronyms List A). To achieve this, We employ a clustering-based method [Li et al., 2023] for partitioning, where a language encoder first generates embeddings, which are then clustered using K-means. Finally, a Dirac distribution with $\alpha=0.5$ is applied to create a label-skewed partitioning [Guo et al., 2024], ensuring each client receives a comparable number of samples.
H.2 Hardware.
All experiments are conducted on a single Nvidia A6000 GPU with 48GB of memory, with a float precision of bf16.
H.3 Experiment Procedure
H.3.1 Federated Utility Fine-Tuning
We begin by performing federated fine-tuning of the LLM[Zhang et al., 2023, Wu et al., 2025] on the partitioned dataset, adapting it to the legal tasks. The fine-tuning is conducted using the OpenFedLLM framework Ye et al. [2024b]. We set the total number of FL rounds to 10 and use FedAVG as the aggregation algorithm.
Each client performs multi-task fine-tuning by mixing all local tasks and applying a unified prompt template, as illustrated in Figure 6, following the approach in Raffel et al. [2023]. In each round of federated learning, the client fine-tunes the received global model for one epoch using parameter-efficient fine-tuning (PEFT) techniques of LoRA [Hu et al., 2021]. The learning rate is set to 3e-4 with a linear decay schedule. The maximum input sequence length is 3072 tokens. We use a batch size of 1 and apply gradient accumulation with a factor of 8. The LoRA configuration is set to $r=16$ and $\alpha=32$.
After federated fine-tuning is complete, we evaluate the utility performance of the final global model on a held-out global test set. In line with standard practices in federated learning research, we also compare this performance with that of a centrally (non-FL) trained model on the same test set. The results are summarized in Table 8.
H.3.2 PII Extraction
In the main experiments, we designate client 0 as the attacker and client 1 as the victim. We construct the prefix set $P_{c}$ for PII-contextual prefix sampling from the local dataset $D_{0}$. During this construction, we set the length parameter $\lambda$ to 50. Each prefix is used to independently query the utility fine-tuned global model $n=15$ times. For each query, the model is allowed to generate up to $m=10$ new tokens. This generation length is sufficient to cover most labeled PII instances while keeping the computational cost acceptable.
For Frequency-Prioritized Prefix Sampling, we construct $\text{Set}(\text{SUP}(P_{c}))$ from the aforementioned $P_{c}$, and sort it in descending order of prefix frequency (as described in Section 4.2.1). The model $\theta$ is then queried using prefixes in this frequency-descending order. Although we do not explicitly define a frequency threshold $\sigma_{a}$, we sweep the prefix budget $B$ exponentially in base 10. Because $\text{Set}(\text{SUP}(P_{c}))$ is sorted by decreasing frequency, this sweep over $B$ implicitly corresponds to sweeping $\sigma_{a}$ from +$\infty$+ ∞ to 1.
H.3.3 Latent Association Fine-tuning
To construct the fine-tuning dataset $D_{\text{ft}}$, we select the top 10000 most frequent prefixes from $\text{Set}(\text{SUP}(P_{c}))$ and randomly sample 10000 PII instances from the attacker’s (client 0’s) PII set $S_{a}$. Although alternative strategies could be explored for prefix and PII selection, this approach is relatively straightforward and effective. We then fine-tune the model $\theta$ to obtain $\theta^{\prime}$ using one epoch and a small learning rate of 5e-5. LoRA is applied with $r=16$ and $\alpha=32$, consistent with the initial federated fine-tuning setup.
Appendix I Resutls of FedLLM Utility Fine-tuning
Below is a task related to judicial and legal matters. Output an appropriately completed response to the request.
<### Input>
{{Task Input}}
<### Output>
{{Task Output}}
Figure 6: Unified Utility Fine-tuning Template for All Tasks.
Table 8: Utility Performance over Different Tasks.
FL Algorithms Models SUM(rouge-l) RC(rouge-l) Match(Acc) Exame(Acc) Cls(Acc)
FedAvg Qwen1-8B 50.0 14.2 50.0 37.5 90.0
FedAvg Baichuan2-7B 57.6 42.4 50.0 33.3 89.5
Non-FL Qwen1-8B 50.0 18.9 50.0 40.8 87.0
Appendix J Ablation Study Results
Table 9: Coverage rate (CR) of attacks for different attacker–victim client pairs with a prefix budget $B=10000$. Prefixes are randomly sampled from the attacker’s corresponding set $P_{c}$. "–" denotes cases where the attacker and victim are the same client (not applicable).
Attacker ID Victim ID Average
0 1 2 3 4
0 - 10.91% 12.89% 10.93% 11.88% 11.66%
1 11.97% - 12.41% 11.46% 12.35% 12.05%
2 12.56% 11.39% - 11.65% 12.74% 12.09%
3 12.07% 10.82% 12.04% - 11.99% 11.73%
4 12.26% 11.36% 13.25% 11.21% - 12.02%
Figure 7: Document frequency distribution of the 2017 VxPII instances extracted from the model trained on the masked dataset.
Figure 8: Label distribution of FedLLM-exclusive VxPII extracted using prefix set $P_{c}$ and the Qwen1-8B model.